OpenStudy (anonymous):

Gaussian Elimination answer check please! 2x+y-z=0 I got x=56/3, y=-20/3, z=8/3 x-y+2z=-3 -x+2y-3z=7

hartnn (hartnn):

there is some error in steps, you need to show your steps, i'll help you spot it and correct it.

OpenStudy (anonymous):

ok hold on i'm drawing it

hartnn (hartnn):

how did you get -2 ? if you have done R2=R2+R3 then 2-3 =-1 and not -2

OpenStudy (anonymous):

which step is that? I went downwards direction

hartnn (hartnn):

oh, you did R3=R2+R3 ? its very unclear to read...

OpenStudy (anonymous):

sorry. here's a better picture if what i did was unclear